# A Progressive Image Restoration Network for High-order Degradation Imaging in Remote Sensing

### v1

@@ Abstract
Recently, deep learning methods have gained remarkable achievements in the field of image restoration for remote sensing (RS). However, most existing RS image restoration methods focus mainly on conventional first-order degradation models, which may not effectively capture the imaging mechanisms of remote sensing images. Furthermore, many RS image restoration approaches that use deep learning are often criticized for their lacks of architecture transparency and model interpretability. To address these problems, we propose a novel progressive restoration network for high-order degradation imaging (HDI-PRNet), to progressively restore different image degradation. HDI-PRNet is developed based on the theoretical framework of degradation imaging, offering the benefit of mathematical interpretability within the unfolding network. The framework is composed of three main components: a module for image denoising that relies on proximal mapping prior learning, a module for image deblurring that integrates Neumann series expansion with dual-domain degradation learning, and a module for super-resolution. Extensive experiments demonstrate that our method achieves superior performance on both synthetic and real remote sensing images.
@@
High-order degradation model, image restoration, remote sensing image, deep unfolding network.
@@ I Introduction
The restoration of degradation images in remote sensing is a challenging task, as it involves restoring the sharp HR image $\bm{u}$ from the deteriorated LR image $\bm{g}$. This problem (1) is considered ill-posed because there are multiple HR solutions that can map to the same LR image, resulting in a non-unique solution. Various methods [2, 3, 4] for image restoration have been investigated. Although many algorithms ignore blur and noise degradation in the forward model (1) and rely solely on the ideal bicubic operator to simulate the super-resolution degradation model, these approaches do not align with the actual physical imaging mechanism of remote sensing, rendering these algorithms impractical in real-world remote sensing applications [5, 6].
Recently, numerous deep learning-based methods have been proposed for image restoration [7, 8, 9, 10, 11, 12, 13, 14] and have demonstrated strong performance across various datasets and evaluation metrics. These methods aimed to integrate advanced deep learning modules and techniques such as channel attention mechanisms [10], spatial attention mechanisms [11], transformer mechanisms [12], and generative adversarial networks (GANs) [15]. The primary aim of these methods is to learn the mapping between LR and HR images. However, several challenges remain. 1) Most super-resolution algorithms use the first-order degradation model, which is inadequate for real images as these are often affected by high-order complex degradations; 2) the end-to-end black-box nature of deep learning-based algorithms makes it challenging to distinguish how the network architecture influences its prediction, emphasizing the need for enhanced transparency and interpretability in these frameworks; 3) many deep learning techniques focus on designing sophisticated network modules and depend on expanding them deeper and wider to enhance performance, this results in increased computational needs and challenges related to the smart design of network architectures.
@@
  • We introduce a novel high-order progressive remote sensing restoration network, termed HDI-PRNet, constructed around a high-order backward degradation model. Our framework adeptly manages complex degradation mechanisms, notably enhancing restoration performance, particularly in real remote sensing applications. The structures aligning backward degradation tasks with respective modules render HDI-PRNet more transparent and interpretable.
@@ III-A High-order Degradation Model
where $\mathcal{D}_{\ell}$ denotes the $\ell$-th imaging sub-system or degradation sensor ($1\leq\ell\leq k$).
In this work, we decompose the $k$-order degradation framework into a series of $k$ sequential first-order models, described as follows
$\begin{split}\bm{g}_{1}=&\mathcal{D}_{1}(\bm{u})=\mathcal{N}_{1}(\mathcal{S}_{1}(\mathcal{B}_{1}(\bm{u})))\\ \bm{g}_{2}=&\mathcal{D}_{2}(\bm{g}_{1})=\mathcal{N}_{2}(\mathcal{S}_{2}(\mathcal{B}_{2}(\bm{g}_{1})))\\ &\cdots\\ \bm{g}:=&\bm{g}_{k}=\mathcal{D}_{k}(\bm{g}_{k-1})=\mathcal{N}_{k}(\mathcal{S}_{k}(\mathcal{B}_{k}(\bm{g}_{k-1}))),\end{split}$ (3)
where the functions $\mathcal{B}_{k}(\cdot)$, $\mathcal{S}_{k}(\cdot)$, $\mathcal{N}_{k}(\cdot)$ represent the blur, downsampling, and noise degradation operators, respectively. The formula (3) outlines a hierarchical degradation model that provides a more precise representation of the real remote sensing imaging system.
III-B Algorithm
The restoration algorithm based on Eq. (3) can be divided into $k$ stages to be solved progressively, where $k$ stages correspond to the high-order degradation mechanism. We assume $\bm{g}_{0}=\bm{u}$, $\bm{g}_{\ell}=\mathcal{D}_{\ell}(\bm{g}_{\ell-1})$ ($1\leq\ell\leq k$) and $\bm{g}_{k}=\bm{g}$. The energy minimization model at $\ell$-stage can be written as
@@ III-B Algorithm
where the regularization $f_{\ell}(\cdot)$ can be used to constrain the geometric prior of a given image in denoising subproblem.
There have been several excellent explorations in solving (4). One of main categories involves step-by-step degradation representations and typically consists of three components: prior denoising or local filtering, super-resolution or interpolation, and blind degradation prediction. The above three components are performed either separately (iteratively) or jointly. Therefore, the above energy minimization model (4) can be divided into three sub-problems, as follows
@@
III-C Proposed Framework
III-C1 Overall Architecture
@@
III-C2 Image denoising proximal mapping module
@@
III-C3 Image super-resolution module
@@
III-C4 Image deblurring module
@@
to learn (10), as shown in Fig. 1(d) and (h). Here, we split the blur operator $\mathcal{B}_{\ell}$ into a multi-scale dual-domain feature representation. Therefore, DDLB includes the spatial transformation branch $\mathcal{B}^{Spatial}_{\ell}$, the frequency transformation branch $\mathcal{B}^{Frequency}_{\ell}$, the dual-domain expression aggregation (DEA) and the identity transformation $\mathcal{I}$.
@@ III-D Loss Function
III-D1 Image reconstruction loss
@@ IV-A1 Datasets
In our experimental investigation of different restoration methods for remote sensing, we have chosen to focus on well-known datasets, namely AID [40], WHU-RS19 [41], and WHU Building [42].
@@
For training and assessment of our methodology, the AID and WHU Building sets are used for training, while WHU-RS19 serves as a test set.
@@ IV-A3 Evaluation Metrics
Figure 3: Visualization comparison for super-resolution scales x2, x3 and x4 on WHU-RS19 dataset.
@@ IV-A4 Compared Methods
We conduct a group of studies with fourteen other methods to evaluate the performance of the proposed method. We compare our MVMS-RCN with general image restoration approaches such as RCAN [10], SwinIR [12], RDN [17], SAN [20], DAT [22], D-DBPN [45], CAT [46], and RGT [47], as well as super-resolution methods in remote sensing, such as MHAN [24], TTST [26], HAUNet [27], TransENet [28], CTNet [48], and HSENet [49].
@@ IV-A5 Training Settings
IV-B Ablation Study
IV-B1 Evaluation for high-order degradation
To demonstrate the advantages of high-order framework, we perform ablation studies, as shown in Table I. Our analysis assesses the performance of framework with differing degradation orders across various degraded images to demonstrate their robustness. We can see from Table I that the low-order degradation network exhibits significantly decreasing performance when faced with complex high-order degraded images. In contrast, high-order degradation network achieves more consistent and superior results in different degradation images, underscoring their increased robustness. For optimal configuration, we chose the second-order degradation architecture because it provides a suitable balance between network complexity and reconstruction performance.
IV-B2 Sensitivity of the loss weight $\alpha$
The WHU-RS19 dataset is used to explore the optimal value for the loss weight $\alpha$. As shown in Table II, setting the weight $\alpha=0$ results in inferior image reconstruction performance compared to the methods where $\alpha\neq 0$. This indicates that the intermediate loss that supervises each submodule is advantageous.
IV-B3 Denoising module architecture

### v2

@@ Abstract
Recently, deep learning methods have gained remarkable achievements in the field of image restoration for remote sensing (RS). However, most existing RS image restoration methods focus mainly on conventional first-order degradation models, which may not effectively capture the imaging mechanisms of remote sensing images. Furthermore, many RS image restoration approaches that use deep learning are often criticized for their lacks of architecture transparency and model interpretability. To address these problems, we propose a novel progressive restoration network for high-order degradation imaging (HDI-PRNet), to progressively restore different image degradation. HDI-PRNet is developed based on the theoretical framework of degradation imaging, also Markov properties of the high-order degradation process and Maximum a posteriori (MAP) estimation, offering the benefit of mathematical interpretability within the unfolding network. The framework is composed of three main components: a module for image denoising that relies on proximal mapping prior learning, a module for image deblurring that integrates Neumann series expansion with dual-domain degradation learning, and a module for super-resolution. Extensive experiments demonstrate that our method achieves superior performance on both synthetic and real remote sensing images.
@@
High-order degradation model, image restoration, Markov process, remote sensing image, deep unfolding network.
@@ I Introduction
The restoration of degradation images in remote sensing is a challenging task, as it involves restoring the sharp HR image $\bm{u}$ from the deteriorated LR image $\bm{g}$. This problem (1) is considered ill-posed because there are multiple HR solutions that can map to the same LR image, resulting in a non-unique solution. Various methods [2, 3, 4, 5, 6, 7] for image restoration have been investigated. Although many algorithms ignore blur and noise degradation in the forward model (1) and rely solely on the ideal bicubic operator to simulate the super-resolution degradation model, these approaches do not align with the actual physical imaging mechanism of remote sensing, rendering these algorithms impractical in real-world remote sensing applications [8, 9].
Recently, numerous deep learning-based methods have been proposed for image restoration [10, 11, 12, 13, 14, 15, 16, 17, 18, 19] and have demonstrated strong performance across various datasets and evaluation metrics. These methods aimed to integrate advanced deep learning modules and techniques such as channel attention mechanisms [14], spatial attention mechanisms [16], transformer mechanisms [17], and generative adversarial networks (GANs) [20]. The primary aim of these methods is to learn the mapping between LR and HR images. However, several challenges remain. 1) Most super-resolution algorithms use the first-order degradation model, which is inadequate for real images as these are often affected by high-order complex degradations; 2) the end-to-end black-box nature of deep learning-based algorithms makes it challenging to distinguish how the network architecture influences its prediction, emphasizing the need for enhanced transparency and interpretability in these frameworks; 3) many deep learning techniques focus on designing sophisticated network modules and depend on expanding them deeper and wider to enhance performance, this results in increased computational needs and challenges related to the smart design of network architectures.
@@
  • We introduce a novel high-order progressive remote sensing restoration network, termed HDI-PRNet, constructed by a reverse Markov process. Our framework adeptly manages complex degradation mechanisms as a Markov chain, notably enhancing restoration performance, particularly in real remote sensing applications. The structures aligning reverse degradation tasks with respective modules render HDI-PRNet more transparent and interpretable.
@@ III-A High-order Degradation Model
where $\mathcal{D}_{\ell}$ denotes the $\ell$-th imaging degradation ($1\leq\ell\leq k$).
In this work, the formula (3) can be modeled as a diffusion-based forward process. We assume $\bm{g}_{0}=\bm{u}$, $\bm{g}_{\ell}=\mathcal{D}_{\ell}(\bm{g}_{\ell-1})$ ($1\leq\ell\leq k$) and $\bm{g}_{k}=\bm{g}$, then decompose the $k$-order degradation framework into a series of $k$ sequential first-order models, described as follows.
$\begin{split}\bm{g}_{1}=&\mathcal{D}_{1}(\bm{g}_{0})=\mathcal{N}_{1}(\mathcal{S}_{1}(\mathcal{B}_{1}(\bm{g}_{0}))),\\ \bm{g}_{2}=&\mathcal{D}_{2}(\bm{g}_{1})=\mathcal{N}_{2}(\mathcal{S}_{2}(\mathcal{B}_{2}(\bm{g}_{1}))),\\ &\cdots\\ \bm{g}:=&\bm{g}_{k}=\mathcal{D}_{k}(\bm{g}_{k-1})=\mathcal{N}_{k}(\mathcal{S}_{k}(\mathcal{B}_{k}(\bm{g}_{k-1}))),\end{split}$ (4)
where $\mathcal{B}_{k}(\cdot)$, $\mathcal{S}_{k}(\cdot)$, $\mathcal{N}_{k}(\cdot)$ represent the blur, downsampling, and noise degradation operators, respectively.
III-B Markov-Chain and MAP Estimation
We formulate the higher-order degradation model (3) as a Markov chain (see Fig.1(a)), so the joint probability distribution of the degraded observations $\bm{g}_{1:k}$ conditioned on the initial clean image $\bm{g}_{0}$ can be expressed as
$q\left(\bm{g}_{1:k}\mid\bm{g}_{0}\right)=\prod_{\ell=1}^{k}q\left(\bm{g}_{\ell}\mid\bm{g}_{\ell-1}\right),$ (5)
where each transition $q\left(\bm{g}_{\ell}\mid\bm{g}_{\ell-1}\right)$ corresponds to a degradation step. The formula (5) outlines a hierarchical degradation model that provides a more precise representation of the real remote sensing imaging system.
The corresponding image restoration aims to recover the original image by progressively reversing the degradation. For degraded observation $\bm{g}_{k}$, this reverse Markov process is modeled as
$\begin{split}p_{\theta}\left(\bm{g}_{0:k-1}|\bm{g}_{k}\right)=\prod_{\ell=1}^{k}p_{\theta}\left(\bm{g}_{\ell-1}\mid\bm{g}_{\ell}\right)\end{split}$ (6)
to find the most probable recovery sequence $\bm{g}_{0:k-1}$ through a Maximum a posteriori (MAP) estimation model defined by
$\begin{split}\underset{\bm{g}_{0:k-1}}{\arg\max}\;p_{\theta}\left(\bm{g}_{0:k-1}|\bm{g}_{k}\right)=\underset{\bm{g}_{0:k-1}}{\arg\max}\prod_{\ell=1}^{k}p_{\theta}\left(\bm{g}_{\ell-1}\mid\bm{g}_{\ell}\right),\end{split}$ (7)
which is progressively solved by
$\bm{g}_{\ell-1}=\arg\max\limits_{\bm{g}}p_{\theta}\left(\bm{g}\mid\bm{g}_{\ell}\right),\quad\ell=k,k-1,\dots,1.$ (8)
By applying Bayes’ formula, the reverse diffusion step splits into components
$p_{\theta}(\bm{g}|\bm{g}_{\ell})\propto q_{\theta}(\bm{g}_{\ell}|\bm{g})p_{\theta}(\bm{g}),\quad\ell=k,k-1,\dots,1,$ (9)
and taking the negative log-likelihood refines the optimization (8) as follow
$\bm{g}_{\ell-1}=\underset{\bm{g}}{\arg\max}\{-\log(q_{\theta}(\bm{g}_{\ell}\mid\bm{g}))-\log(p_{\theta}(\bm{g}))\}.$ (10)
Further, we consider the following assumptions
$\begin{split}-\log(q_{\theta}(\bm{g}_{\ell}\mid\bm{g}))&\propto\frac{1}{2\sigma_{\ell}^{2}}\|\bm{g}_{\ell}-\mathcal{S}_{\ell}(\mathcal{B}_{\ell}(\bm{g}))\|,\\ -\log(p(\bm{g}))&\propto\frac{1}{2\delta_{\ell}^{2}}f_{\ell}(\bm{g}).\end{split}$ (11)
Thus, the restoration algorithm based on Eq. (3) and Eq. (4) can be divided into $k$ stages to be solved progressively, where $k$ stages correspond to the high-order degradation mechanism. The energy minimization model (10) at $\ell$-stage can be rewritten as
@@ III-B Markov-Chain and MAP Estimation
where the regularization $f_{\ell}(\cdot)$ can be used to constrain the geometric prior of a given image in the denoising subproblem.
III-C Algorithm
There have been several excellent explorations [40, 41, 39] in solving (12). One of main categories involves step-by-step degradation representations and typically consists of three components: prior denoising or local filtering, super-resolution or interpolation, and blind degradation prediction. The above three components are performed either separately (iteratively) or jointly. Therefore, the above energy minimization model (12) can be divided into three sub-problems, as follows
@@ III-C Algorithm
III-D Proposed Framework
III-D1 Overall Architecture
@@
III-D2 Image denoising module
@@
III-D3 Image super-resolution module
@@
III-D4 Image deblurring module
@@
to learn (19), as shown in Fig. 1(d) and (h). Here, we split the blur operator $\mathcal{B}_{\ell}$ into a dual-domain feature representation. Therefore, DDLB includes the spatial transformation branch $\mathcal{B}^{Spatial}_{\ell}$, the frequency transformation branch $\mathcal{B}^{Frequency}_{\ell}$, the dual-domain expression aggregation (DEA) and the identity transformation $\mathcal{I}$.
@@ III-E Loss Function
III-E1 Image reconstruction loss
@@ IV-A1 Datasets
In our experimental investigation of different remote sensing restoration methods, we have chosen to focus on well-known datasets, namely AID [46], WHU-RS19 [47], and WHU Building [48].
@@
  • 4)
    DOTA Dataset [49] comes from a variety of different sensors, including Google Earth, Jilin-1 satellite and Gaofen-2 satellite. The size of each image ranges from $800\times 800$ to $4000\times 4000$.
  • 5)
    RSSCN7 Dataset [50] has 2,800 images, including 7 classic ground feature categories: Grass, Field, Industry, River Lake, Forest, Resident, Parking. Each category contains 400 images, and the images come from different seasons and weather conditions.
  • 6)
    UCMerced Dataset [51] is a remote sensing image dataset containing 21 types of land objects, each category contains 100 images.
  • 7)
    NWPU-RESISC45 Dataset [52] contains 45 types of remote sensing scene images with a size of $256\times 256$, covering a variety of scene categories such as airports, beaches, bridges, etc.
For training and assessment of our methodology, the AID and WHU Building sets are used for training, while other datasets serve as testing.
@@ IV-A3 Evaluation Metrics
Figure 3: $\times$2, $\times$3 and $\times$4 visual comparisons on DOTA, NWPU-RESISC45, UCMerced, RSSCN7 and WHU-RS19 datasets.
@@ IV-A4 Compared Methods
In this study, we conduct a group of studies with fourteen methods to evaluate the performance of the proposed method. Specifically, we compare our HDI-PRNet with general image restoration approaches such as RCAN [14], SwinIR [17], RDN [22], SAN [25], DAT [27], D-DBPN [55], CAT [56], and RGT [57], as well as super-resolution methods in remote sensing, such as MHAN [29], TTST [31], HAUNet [32], TransENet [33], CTNet [58], and HSENet [59]. To ensure consistency and rigor, all methods included in the study were implemented using the original code provided by the respective authors, available on their official homepages or GitHub repositories, and re-trained on the same training dataset.
@@ IV-A5 Training Settings
IV-B Ablation Study
IV-B1 Evaluation for high-order degradation
To demonstrate the advantages of the high-order framework, we perform ablation studies, as shown in Table I. Our analysis assesses the performance of frameworks with differing degradation orders across various degraded images to demonstrate their robustness. We can see from Table I that the low-order degradation network exhibits significantly decreasing performance when faced with complex high-order degraded images. In contrast, a high-order degradation network achieves more consistent and superior results in different degradation images, underscoring their increased robustness. For optimal configuration, we chose the second-order degradation architecture because it provides a suitable balance between network complexity and reconstruction performance.
IV-B2 Sensitivity of the loss weight $\alpha$
The WHU-RS19 dataset is used to explore the optimal value of loss of weight $\alpha$. As shown in Table II, setting the weight $\alpha=0$ results in inferior image reconstruction performance compared to the methods where $\alpha\neq 0$. This indicates that the intermediate loss that supervises each submodule is advantageous.
IV-B3 Denoising module architecture
@@ IV-B4 Deblurring module architecture
It should be noted that our Neumann first-order expansion network, even in its most basic form, outperforms both DAT and TTST in reconstruction performance.
TABLE VI: Comparison of different methods on various remote sensing datasets. The best and second-best results are highlighted in Bold font and the underlined ones, respectively.
Dataset Scale WHU-RS19 DOTA RSSCN7 UCMerced NWPU-RESISC45
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
CTNet [58] ×$\times 2$ 28.64 0.7569 30.09 0.7876 27.39 0.6799 26.43 0.7048 27.02 0.6975
HSENet [59] ×$\times 2$ 28.94 0.7616 30.70 0.7964 27.71 0.6853 26.79 0.7114 27.40 0.7033
D-DBPN [55] ×$\times 2$ 29.20 0.7698 30.41 0.7987 27.81 0.6920 26.99 0.7231 27.52 0.7122
RDN [22] ×$\times 2$ 29.33 0.7740 30.85 0.8053 27.88 0.6956 27.06 0.7290 27.63 0.7159
TransENet [33] ×$\times 2$ 29.33 0.7756 30.92 0.8079 27.89 0.6977 27.13 0.7308 27.66 0.7177
RCAN [14] ×$\times 2$ 29.46 0.7786 31.03 0.8113 27.98 0.7016 27.23 0.7350 27.72 0.7214
SAN [25] ×$\times 2$ 29.38 0.7742 30.99 0.8075 27.94 0.6970 27.14 0.7290 27.68 0.7169
MHAN [29] ×$\times 2$ 29.40 0.7761 30.87 0.8076 27.92 0.6976 27.14 0.7305 27.68 0.7184
SwinIR [17] ×$\times 2$ 29.41 0.7778 30.97 0.8100 27.93 0.7008 27.16 0.7352 27.68 0.7206
HAUNet [32] ×$\times 2$ 29.52 0.7789 31.25 0.8131 28.01 0.7003 27.37 0.7399 27.78 0.7205
TTST [31] ×$\times 2$ 29.45 0.7771 31.09 0.8118 27.97 0.7008 27.25 0.7356 27.73 0.7200
CAT [56] ×$\times 2$ 29.63 0.7838 31.12 0.8152 28.05 0.7053 27.39 0.7427 27.82 0.7253
RGT [57] ×$\times 2$ 29.63 0.7840 31.17 0.8156 28.09 0.7058 27.39 0.7430 27.86 0.7263
DAT [27] ×$\times 2$ 29.68 0.7855 31.22 0.8166 28.11 0.7074 27.42 0.7448 27.89 0.7276
HDI-PRNet ×$\times 2$ 30.03 0.7931 31.66 0.8230 28.32 0.7135 27.82 0.7542 28.13 0.7334
CTNet [58] ×$\times 3$ 27.59 0.7200 29.39 0.7603 26.72 0.6428 25.61 0.6652 26.32 0.6579
HSENet [59] ×$\times 3$ 27.98 0.7292 29.74 0.7691 26.98 0.6500 25.86 0.6756 26.50 0.6656
D-DBPN [55] ×$\times 3$ 28.31 0.7431 29.76 0.7767 27.20 0.6616 26.26 0.6927 26.79 0.6798
RDN [22] ×$\times 3$ 28.38 0.7449 30.05 0.7807 27.21 0.6622 26.20 0.6938 26.79 0.6803
TransENet [33] ×$\times 3$ 28.44 0.7482 30.05 0.7828 27.24 0.6642 26.27 0.6971 26.81 0.6818
RCAN [14] ×$\times 3$ 28.48 0.7513 30.17 0.7865 27.27 0.6681 26.29 0.6991 26.85 0.6852
SAN [25] ×$\times 3$ 28.31 0.7435 30.03 0.7801 27.19 0.6616 26.19 0.6927 26.76 0.6789
MHAN [29] ×$\times 3$ 28.45 0.7485 30.10 0.7836 27.26 0.6654 26.30 0.6986 26.87 0.6842
SwinIR [17] ×$\times 3$ 28.53 0.7523 30.15 0.7855 27.29 0.6686 26.31 0.7015 26.88 0.6854
HAUNet [32] ×$\times 3$ 28.51 0.7483 30.25 0.7855 27.31 0.6649 26.44 0.7045 26.96 0.6861
TTST [31] ×$\times 3$ 28.51 0.7528 30.32 0.7887 27.32 0.6693 26.41 0.7036 26.90 0.6864
CAT [56] ×$\times 3$ 28.67 0.7573 30.39 0.7918 27.38 0.6731 26.45 0.7084 26.97 0.6904
RGT [57] ×$\times 3$ 28.70 0.7580 30.39 0.7919 27.42 0.6742 26.47 0.7083 27.00 0.6908
DAT [27] ×$\times 3$ 28.72 0.7587 30.37 0.7925 27.41 0.6748 26.48 0.7088 27.01 0.6918
HDI-PRNet ×$\times 3$ 28.96 0.7623 30.72 0.7956 27.57 0.6766 26.85 0.7157 27.22 0.6953
CTNet [58] ×$\times 4$ 26.74 0.6707 28.68 0.7274 26.20 0.6040 24.78 0.6175 25.68 0.6182
HSENet [59] ×$\times 4$ 27.49 0.7066 29.29 0.7531 26.64 0.6323 25.46 0.6560 26.15 0.6465
D-DBPN [55] ×$\times 4$ 27.42 0.7049 29.23 0.7508 26.63 0.6311 25.46 0.6528 26.15 0.6457
RDN [22] ×$\times 4$ 27.42 0.7050 29.29 0.7503 26.61 0.6308 25.41 0.6537 26.15 0.6455
TransENet [33] ×$\times 4$ 27.53 0.7102 29.40 0.7568 26.66 0.6354 25.54 0.6612 26.20 0.6500
RCAN [14] ×$\times 4$ 27.55 0.7127 29.34 0.7570 26.68 0.6372 25.52 0.6629 26.20 0.6523
SAN [25] ×$\times 4$ 27.38 0.7024 29.28 0.7494 26.58 0.6291 25.36 0.6510 26.09 0.6424
MHAN [29] ×$\times 4$ 27.48 0.7071 29.38 0.7543 26.63 0.6327 25.44 0.6566 26.17 0.6475
SwinIR [17] ×$\times 4$ 27.65 0.7152 29.39 0.7579 26.72 0.6391 25.60 0.6657 26.25 0.6533
HAUNet [32] ×$\times 4$ 27.59 0.7117 29.47 0.7586 26.71 0.6362 25.68 0.6669 26.27 0.6521
TTST [31] ×$\times 4$ 27.61 0.7138 29.40 0.7587 26.70 0.6376 25.57 0.6639 26.23 0.6521
CAT [56] ×$\times 4$ 27.70 0.7176 29.56 0.7618 26.76 0.6407 25.67 0.6692 26.28 0.6553
RGT [57] ×$\times 4$ 27.79 0.7199 29.59 0.7644 26.80 0.6430 25.72 0.6729 26.33 0.6579
DAT [27] ×$\times 4$ 27.79 0.7203 29.59 0.7639 26.80 0.6431 25.75 0.6721 26.35 0.6579
HDI-PRNet ×$\times 4$ 27.87 0.7214 29.78 0.7655 26.91 0.6434 25.92 0.6732 26.46 0.6586